if 17 is the number of sequence and the next number is 3 what is the first number before 17 thanks
Answer by Fombitz(32388)

I assume you mean you have a sequence of three numbers, the first is X, the second is 17 and the third is 3.
The difference between 3 and 17 is the same as the difference between 17 and X.
3-17=17-X
X=34-3
X=31
31,17,3